OpenStudy (mendicant_bias):

I have a vaguely ODE-related trigonometry problem dealing with the Differential Equation for Simple Harmonic Motion, question posted below momentarily.

OpenStudy (mendicant_bias):

I was looking at how to solve the ODE \[F = m \frac{d^2x}{dt^2}=-kx,\]or\[\frac{d^2x}{dt^2}+kx=0\] resulting from equating Hooke's Law and Newton's 2nd Law, is solved, and I understanding how the roots are obtained for the char. Eqn and all, giving you \[c_{1}\cos(\omega t)+c_{2}\sin(\omega t),\]where omega is equal to sqrt(k/m).

OpenStudy (mendicant_bias):

What I don't get, though, is that they are then equated to, the general equation for SHM, something like \[c_{1}\cos(\omega t)+c_{2}\sin(\omega t)=A sin (\omega t + \zeta)\]

OpenStudy (mendicant_bias):

(I used the wrong Greek letter, but I couldn't remember what the other one was in the argument of sine, lol) But yeah, is there a trig identity being employed or something there? It's on the Wiki page for SHM

OpenStudy (mendicant_bias):

This is exactly the equality I'm talking about: http://upload.wikimedia.org/math/6/5/6/656fd81e91b7ad38db0c1f263dd5f4af.png

OpenStudy (anonymous):

\[\sin(\alpha\pm\beta)=\sin\alpha\cos\beta\pm\sin\beta\cos\alpha\] where \(\alpha=\omega t\), \(\beta=\zeta\), \(\cos\beta=c_1\) and \(\sin\beta=c_2\).

OpenStudy (anonymous):

(and the Greek letter is phi, or `\varphi` in latex, giving \(\varphi\))

OpenStudy (mendicant_bias):

Oh, wow, that's a strange way of employing *that trig identity; I've never seen it done like that before. Neat.

OpenStudy (mendicant_bias):

That's still very strange, I'm trying to wrap my head around how the arguments of the sine functions equal to the constants would work. Just going to think about it for a little.

OpenStudy (anonymous):

Above I suppose I'm fixing \(A=1\), but in general you would have \(c_1=A\cos\zeta\) and \(c_2=A\sin\zeta\).